OpenStudy (anonymous):

I have a question asking what is equal to -sec^2x. I believe that the answer is -tan^2x-1. However, I am not completely positive

OpenStudy (anonymous):

I came to this conclusion by making 1+ tan^2 =sec^2 all negative to get -1 - tan^2x = -sec^2 which is also equal to -tan^2x -1 = -sec^2x. Is this correct or have i done something wrong? Please helpp

OpenStudy (amistre64):

its correct

OpenStudy (anonymous):

okay thank you!

OpenStudy (amistre64):

\[sec^2=\frac1{cos^2}\] would also seem fine

OpenStudy (anonymous):

well it was a multiple choice questions with the answers: a. tan^2 x -1 b. -tan^2 x +1 c. -cos^2 x d. -tan^2 x -1 e. none of the above I said that it was D. @amistre64

OpenStudy (amistre64):

D does fit best with the options given :)

OpenStudy (anonymous):

ok thanks :)